• Given the demand for a product as Qd = 200 - 2P and the supply is given as Qs = 50+P. You are told equilibrium is obtained at the point where Qd = Qs. The equilibrium quantity for the product is A. 200
  •  B. 250
  •  C. 100
  •  D. 150
1
Expert's answer
2021-06-21T11:57:32-0400

Let's first find the equilibrium price:


"Q_d=Q_s,""200 - 2P_E=50+P_E,""150=3P_E,""P_E=R50."


Finally, we can find the equilibrium quantity:


"Q_E=200 - 2\cdot R50=100."

Therefore, the correct answer is (C) 100.